# Cost-Effective Text Clustering with Large Language Models

### v1

@@ 7.4. Ablation Study and Component Analysis
Figure 6. The performance of TECL with various LLMs
@@
Queried LLMs. Fig. 6 illustrates the clustering results of TECL when querying 4 popular LLMs: GPT-4o, GPT-4o-Mini, Gemini-1.5-Pro, and Claude-3.5-Sonnet. Key observations include: (1) Minimal differences in the adoption of various LLMs. (2) LLMs influence answer accuracy, particularly regarding constraints, without affecting query triangle selection; for example, GPT-4o enhances constraint accuracy compared to GPT-4o-Mini, resulting in increased overall accuracy. (3) The advantages of more expensive LLMs do not warrant their costs, underscoring the necessity for more cost-effective models capable of managing complex tasks like triangle queries.
8. Conclusion
This paper presents TECL, a framework that cost-effectively leverages LLMs to produce accurate pair-wise constraints for subsequent constrained text clustering. In particular, we develop two greedy algorithms EdgeLLM and TriangleLLM to carefully select text pairs and triplets for querying LLMs so as to obtain the most informative constraints for clustering, and further ameliorate standard constrained clustering algorithms with personalized weights for various pair constraints. Our extensive experiments over benchmark datasets exhibit the high superiority of our proposed techniques in enhancing text clustering quality and reducing the sheer amount of queries needed for LLMs.

### v2

@@ Appendix D Additional Experiments
Figure 6. The performance of Cequel with various LLMs
@@
Queried LLMs. Fig. 6 illustrates the clustering results of Cequel when querying four popular LLMs including GPT-4o, GPT-4o-Mini, Gemini-1.5-Pro, and Claude-3.5-Sonnet. Key observations include: (1) Minimal differences in the adoption of various LLMs. (2) LLMs influence answer accuracy, particularly regarding constraints, without affecting query triangle selection; for example, GPT-4o enhances constraint accuracy compared to GPT-4o-Mini, resulting in increased overall accuracy. (3) The advantages of more expensive LLMs do not warrant their costs, underscoring the necessity for more cost-effective models capable of managing complex tasks like triangle queries.
Appendix E Additional Prompts
Table 8. Example prompts for edge and triangle queries on different datasets.
Dataset Type Prompt
BBC News Edge Query Cluster BBC News docs by whether they belong to the same news category. For each pair, respond with Yes or No without explanation.
- News #1: Ad sales boost Time Warner profit
- News #2: Air passengers win new EU rights
Given this context, do News #1 and News #2 likely correspond to the same news category?
Triangle Query Cluster BBC News docs by whether they belong to the same news category. For each triangle, respond with a, b, c, d, or e without explanation.
- News #1: REM announce new Glasgow concert
- News #2: Moreno debut makes Oscar mark
- News #3: Last Star Wars ’not for children’
Given this context, do News #1, News #2, and News #3 likely correspond to the same news category?
(a) All are same category. (b) Only #1 and #2 are same category. (c) Only #1 and #3 are same category.
(d) Only #2 and #3 are same category. (e) None.
Tweet Edge Query Cluster Tweet docs by whether they belong to the same tweet category. For each pair, respond with Yes or No without explanation.
- Tweet #1: super bowl commercial
- Tweet #2: kung pao chicken recipe
Given this context, do Tweet #1 and Tweet #2 likely correspond to the same tweet category?
Triangle Query Cluster Tweet docs by whether they belong to the same tweet category. For each triangle, respond with a, b, c, d, or e without explanation.
- Tweet #1: weight loss diet plan acai juice equal nutritional claim acai
- Tweet #2: watch christina aguilera screw national anthem super bowl post
- Tweet #3: yell cast jr hartley ad digital era author search fly fishing book dj hunt book read
Given this context, do Tweet #1, Tweet #2, and Tweet #3 likely correspond to the same tweet category?
(a) All are same category. (b) Only #1 and #2 are same category. (c) Only #1 and #3 are same category.
(d) Only #2 and #3 are same category. (e) None.
Bank77 Edge Query Cluster Bank77 docs by whether they belong to the same intent category. For each pair, respond with Yes or No without explanation.
- Intent #1: I think something went wrong with my card delivery as I haven’t received it yet.
- Intent #2: How do I link to my credit card with you?
Given this context, do Intent #1 and Intent #2 likely correspond to the same intent category?
Triangle Query Cluster Bank77 docs by whether they belong to the same intent category. For each triangle, respond with a, b, c, d, or e without explanation.
- Intent #1: What is the current exchange rate for me?
- Intent #2: I believe my card payment exchange rate is incorrect.
- Intent #3: I think my child used my card while I wasn’t home.
Given this context, do Intent #1, Intent #2, and Intent #3 likely correspond to the same intent category?
(a) All are same category. (b) Only #1 and #2 are same category. (c) Only #1 and #3 are same category.
(d) Only #2 and #3 are same category. (e) None.
Reddit Edge Query Cluster Reddit docs by whether they belong to the same topic category. For each pair, respond with Yes or No without explanation.
- Topic #1: CTA to make all rail stations accessible within 20 years
- Topic #2: A Krazy Mug Oyee Balle Balle!!! Kettle with set of two glasses
Given this context, do Topic #1 and Topic #2 likely correspond to the same topic category?
Triangle Query Cluster Reddit docs by whether they belong to the same topic category. For each triangle, respond with a, b, c, d, or e without explanation.
- Topic #1: Batman and Batwoman Fan art
- Topic #2: Books on American Union and Labour Movements?
- Topic #3: When Taliban offer you gold: Afghan youth in crisis?
Given this context, do Topic #1, Topic #2, and Topic #3 likely correspond to the same topic category?
(a) All are same category. (b) Only #1 and #2 are same category. (c) Only #1 and #3 are same category.
(d) Only #2 and #3 are same category. (e) None.
Massive Edge Query Cluster Massive docs by whether they belong to the same scenario category. For each pair, respond with Yes or No without explanation.
- Scenario #1: here is something from today
- Scenario #2: remind me of the meeting on tuesday
Given this context, do Scenario #1 and Scenario #2 likely correspond to the same scenario category?
Triangle Query Cluster Massive docs by whether they belong to the same scenario category. For each triangle, respond with a, b, c, d, or e without explanation.
- Scenario #1: alexa play my country playlist
- Scenario #2: write email to my company mate to submit the task tomorrow
- Scenario #3: remind me about my business meeting at three and forty five p. m.
Given this context, do Scenario #1, Scenario #2, and Scenario #3 likely correspond to the same scenario category?
(a) All are same category. (b) Only #1 and #2 are same category. (c) Only #1 and #3 are same category.
(d) Only #2 and #3 are same category. (e) None.
We provide additional example prompts in Table 8 for edge and triangle queries across multiple datasets, including BBC News, Tweet, Bank77, Reddit, and Massive. Edge queries ask whether two documents belong to the same category (Yes/No), while triangle queries assess the categorical relationship among three documents via a five-way multiple-choice format. These prompts are designed to support constraint construction using LLMs for clustering tasks.